Question 13

1 / 1 pts

The following cost data is for the mythical land of Painduvin where they produce nothing but bread and wine using only land and labor as inputs.

1 unit of bread 1 unit of cheese
Labor Input 5 dollars 20 dollars
Land Input 4 dollars 10 dollars


If Painduvin were land-abundant, the opening up of free trade would cause the production of bread to __________ and the production of cheese to __________.

  

Fall; fall

   

Rise; rise

   

Rise; fall

   

Fall; rise
Answer #1

Labor land ratio Of bread= 5/4= 1.25

Labor land ratio of cheese = 20/10= 2

This means cheese is labor intensive and bread is land intensive.

Rise, fall - is correct